OpenStudy (wendy.ivette11714):

x2 + 4x + 3 = 0 Help medal!!

OpenStudy (wendy.ivette11714):

@KendrickLamar2014 @Elsa213 @Luigi0210

OpenStudy (photon336):

look at the last term 3

OpenStudy (igreen):

@ageta Don't give answers

OpenStudy (ageta):

sorry that is not the answer

OpenStudy (wendy.ivette11714):

Let me take a picture

OpenStudy (igreen):

Factor it.

OpenStudy (igreen):

Can you do that?

OpenStudy (photon336):

what two factors of 3 if added together give you 4 the middle term?

OpenStudy (wendy.ivette11714):

http://m.imgur.com/EwJyMLf That's how it looks like

OpenStudy (photon336):

two things to do when factoring. 1. look at the last term your C-term 2. Then determine what two factors of the C term when added give you the B term. 3. be sure to look at the signs also.

OpenStudy (wendy.ivette11714):

I got 191

OpenStudy (wendy.ivette11714):

That's wrong right?

OpenStudy (kendricklamar2014):

To solve this problem, were going to have to factor. We need to look at the constant term (3) and find factors of the contrast term that add up to 4. 3 can be factored only as 3 and 1. \[3 + 1 = 4\]

OpenStudy (kendricklamar2014):

So you can factor in the following way: \[(x+3)(x+1) = 0\] Now we apply the Zero Product Property. This property states that if the whole things is 0, if (x+3)(x+1) is equal to 0, then (x+3) will equal 0, or (x+1) equals 0.

OpenStudy (wendy.ivette11714):

x=-3 or x=-1 ?

OpenStudy (kendricklamar2014):

(x+3) is equal to 0 when x itself is equal to -3 (x = -3). This is because -3 + 3 = 0. Or x is equal to -1 would be equal to 0 because -1 + 1 = 0. So there are two possible answers: \[x = -3 ~ or ~ x= -1\]

OpenStudy (kendricklamar2014):

Yes, that is correct @Wendy.Ivette11714

OpenStudy (wendy.ivette11714):

Okay thank you!! :)

OpenStudy (kendricklamar2014):

Your welcome ^_^